# On the universality of star formation efficiency in galaxies

### v2

@@ 2 Simulations
where the total sub-grid velocity dispersion, $\sigma_{\mathrm{tot}}=\sqrt{\sigma_{\mathrm{t}}^{2}+c_{\mathrm{s}}^{2}}$, includes contributions of both turbulence and the thermal energy of the gas. Note that the virial parameter depends on both the gas temperature through the sound speed and on the turbulent velocity dispersion. This, for example, prevents star formation in the warm gas ($T\gtrsim 10^{3}$ K). Note also that there is no explicit dependence of $\epsilon_{\rm ff}$ on the Mach number because no significant dependence on the Mach number was found in simulations by Padoan et al. (2012) and more recent simulations by Brucy et al. (2024). Figure 1 shows the $\sigma_{\mathrm{tot}}-n$ phase space in our simulations. The dashed line shows the constant value of $\alpha_{\mathrm{vir}}=15$, below which most of the star formation occurs. The figure shows that with the prescription we use in our simulations stars form in cold gas ($T\sim 10-30$ K) with densities of $n\sim 10-1000^{3}\rm\,cm^{-3}$ similar to the densities of observed star-forming molecular clouds averaged on similar scales (tens of parsecs). It also shows that the star-forming gas is different than the gas selected above a constant density and or below a given temperature threshold.
@@ 3.2 Star formation efficiency as a function of local environment properties on 500 pc scale
The efficiency $\epsilon_{\mathrm{ff}}$ shows only a weak trend with surface density. Our results thus indicate that in the simulations with the turbulence-based SFR model with realistic stellar feedback the natural result is the universality of characteristic $\epsilon_{\mathrm{ff}}$ values of star-forming gas in environments with a wide range of metallicities, UV fluxes, and densities and over a wide range of averaging scales. This is further illustrated in Figure 7, which shows the face-on projected maps of $\epsilon_{\mathrm{ff}}$ along with the maps of baryon surface density. The figure visually shows that while surface density varies by more than an order of magnitude in the inner regions of the disk, the values of $\epsilon_{\mathrm{ff}}$ stay in a narrow range around $\epsilon_{\rm ff}\approx 0.01$.
The median free-fall time of star-forming gas, on the other hand, shows tight, nearly metallicity-independent correlations with the UV flux and baryon surface density. This indicates that star-forming gas is denser in denser regions of the galaxy. This may be because gas in high surface density regions experiences a stronger gravitational pull by the disk compressing it to higher densities. The trend with UV flux may reflect the positive correlation of the UV flux and star formation rate, which, in turn, correlates with gas density. These trends of $t_{\mathrm{ff}}$ result in similar trends of the depletion time with $U_{\rm MW}$ and $\Sigma_{\rm\star+g}$.
@@ 4 Discussion
This also allows us to interpret the universality of $\epsilon_{\mathrm{ff}}$ in environments of different metallicity, UV flux, and densities demonstrated above. The pressure and $\alpha_{\rm vir}$ in star-forming regions in these simulations are dominated by turbulence. The turbulent energy is driven by gas compression by large-scale flows and the dissipation on the turbulent crossing time scale. Both processes do not depend on the metallicity, UV flux, or density. In contrast, the gas cooling and heating, and therefore thermal pressure, do depend on these properties and thus, when $\alpha_{\rm vir}$ is dominated by the thermal pressure, the efficiency per free-fall time varies more strongly with the properties of the environment. This is illustrated explicitly in the Appendix A, where we show $\epsilon_{\mathrm{ff}}$ as a function of metallicity, UV flux, and surface density in simulations, in which subgrid turbulence is not modeled and its contribution to $\alpha_{\rm vir}$ is ignored. The turbulence thus plays an important role in making $\epsilon_{\mathrm{ff}}$ value range universal.
@@
The values of $\epsilon_{\mathrm{ff}}$ and their universality exhibited in our simulations are consistent with the typical values of $\approx 0.01-0.1$ estimated for observed gas in diverse environments and averaging scales (e.g., García-Burillo et al., 2012; Utomo et al., 2018; Sun et al., 2023). Estimates of $\epsilon_{\mathrm{ff}}$  from the star-forming sub-pc cores to giant molecular clouds on tens of pc scales are $\approx 0.01-0.1$ (e.g., Krumholz & Tan, 2007; Evans et al., 2014; Pokhrel et al., 2021; Hu et al., 2022). Similar values of $\epsilon_{\mathrm{ff}}$ are estimated on the scales of hundreds of pc in other galaxies with a diverse range of metallicities and surface densities (e.g., Leroy et al., 2008; Utomo et al., 2018; Teng et al., 2024). Our results indicate that these characteristic values of $\epsilon_{\mathrm{ff}}$ and their universality can be plausibly explained by the turbulence-driven and feedback-regulated properties of star-forming regions.
@@
We note, however, that comparing detailed distribution of $\epsilon_{\mathrm{ff}}$ estimated in observations and simulations is not straightforward. In simulations, instantaneous values of $\epsilon_{\mathrm{ff}}$ are considered, while in observations, these values are estimated using instantaneous gas properties and star formation rate indicators that are sensitive to star formation over a certain time scale. The gas density is expected to evolve strongly as the star formation and feedback proceed in a given region (Feldmann & Gnedin, 2011). The large apparent scatter of observational estimates of $\epsilon_{\mathrm{ff}}$ thus largely reflects diverse evolutionary stages of the regions in which these estimates are made (e.g., Feldmann & Gnedin, 2011; Lee et al., 2016; Grisdale et al., 2019). Nevertheless, the simulations shed light on why the characteristic average values of $\epsilon_{\mathrm{ff}}$ do not vary with environment properties.

### v3

@@ 2 Simulations
where the total sub-grid velocity dispersion, $\sigma_{\mathrm{tot}}=\sqrt{\sigma_{\mathrm{t}}^{2}+c_{\mathrm{s}}^{2}}$, includes contributions of both turbulence and the thermal energy of the gas. Note that the virial parameter depends on both the gas temperature through the sound speed and on the turbulent velocity dispersion. This, for example, prevents star formation in the warm gas ($T\gtrsim 10^{3}$ K). Note also that there is no explicit dependence of $\epsilon_{\rm ff}$ on the Mach number because no significant dependence on the Mach number was found in simulations by Padoan et al. (2012) and more recent simulations by Brucy et al. (2024). Figure 1 shows the $\sigma_{\mathrm{tot}}-n$ phase space in our simulations. The dashed line shows the constant value of $\alpha_{\mathrm{vir}}=15$, below which most of the star formation occurs. The figure shows that with the prescription we use in our simulations stars form in cold gas ($T\sim 10-30$ K) with densities of $n\sim 10-10^{3}\rm\,cm^{-3}$ similar to the densities of observed star-forming molecular clouds averaged on similar scales (tens of parsecs). It also shows that the star-forming gas is different than the gas selected above a constant density and or below a given temperature threshold.
@@ 3.2 Star formation efficiency as a function of local environment properties on 500 pc scale
The efficiency $\epsilon_{\mathrm{ff}}$ shows only a weak trend with surface density. Our results thus indicate that in the simulations with the turbulence-based SFR model with realistic stellar feedback, the natural result is the universality of characteristic $\epsilon_{\mathrm{ff}}$ values of star-forming gas in environments with a wide range of metallicities, UV fluxes, and densities and over a wide range of averaging scales. This is further illustrated in Figure 7, which shows the face-on projected maps of $\epsilon_{\mathrm{ff}}$ along with the maps of baryon surface density. The figure visually shows that while surface density varies by more than an order of magnitude in the inner regions of the disk, the values of $\epsilon_{\mathrm{ff}}$ stay in a narrow range around $\epsilon_{\rm ff}\approx 0.01$.  [^3] Note that the trend with surface density shown in Figures 6 and 7 is different from that derived by Sun et al. (2023) for observed galaxies. This is likely because efficiency and free-fall time are defined differently in that study. Namely, the free-fall time is defined as the median time of molecular gas averaged on 60 pc scale, while $\Sigma_{\rm sfr}$ in their study is estimated on the 1.5 kpc scale rather than 500 pc in our case. Moreover, we define both $\epsilon_{\mathrm{ff}}$ and $t_{\rm ff}$ relative to the star-forming gas, which is not the same as molecular gas in our model (see Semenov et al., 2019, for detailed discussion). If our definition $\epsilon_{\rm ff}\approx\Sigma_{\rm sfr}/(\Sigma_{\rm sf}/t_{\rm ff})$, where $\Sigma_{\rm sf}$ is the surface density of star-forming gas, while Sun et al. (2023) define $\epsilon_{\rm ff,mol}=\Sigma_{\rm sfr}/(\Sigma_{\rm mol}/t_{\rm ff,mol})$, where $\Sigma_{\rm mol}$ is the surface density of molecular gas. The two definitions are thus related by $\epsilon_{\rm ff}=(\Sigma_{\rm mol}/\Sigma_{\rm sf})(t_{\rm ff}/t_{\rm ff,mol})\epsilon_{\rm ff,mol}$. The relation depends on the trend of $(\Sigma_{\rm mol}/\Sigma_{\rm sf})(t_{\rm ff}/t_{\rm ff,mol})$ with surface density.
The median free-fall time of star-forming gas, on the other hand, shows tight, nearly metallicity-independent correlations with the UV flux and baryon surface density. This indicates that star-forming gas is denser in denser regions of the galaxy. This may be because gas in high surface density regions experiences a stronger gravitational pull by the disk compressing it to higher densities. The trend with UV flux may reflect the positive correlation between the UV flux and star formation rate, which, in turn, correlates with gas density. These trends of $t_{\mathrm{ff}}$ result in similar trends of the depletion time with $U_{\rm MW}$ and $\Sigma_{\rm\star+g}$.
@@ 4 Discussion
This also allows us to interpret the universality of $\epsilon_{\mathrm{ff}}$ in environments of different metallicity, UV flux, and densities demonstrated above. The pressure and $\alpha_{\rm vir}$ in star-forming regions in these simulations are dominated by turbulence. The turbulent energy is driven by gas compression by large-scale flows and the dissipation on the turbulent crossing time scale. Both processes do not depend on the metallicity, UV flux, or density. In contrast, the gas cooling and heating, and therefore thermal pressure, do depend on these properties and thus, when $\alpha_{\rm vir}$ is dominated by the thermal pressure, the efficiency per free-fall time varies more strongly with the properties of the environment. This is illustrated explicitly in Appendix A, where we show $\epsilon_{\mathrm{ff}}$ as a function of metallicity, UV flux, and surface density in simulations in which subgrid turbulence is not modeled and its contribution to $\alpha_{\rm vir}$ is ignored. The turbulence thus plays an important role in making $\epsilon_{\mathrm{ff}}$ value range universal.
The values of $\epsilon_{\mathrm{ff}}$ and their universality exhibited in our simulations are consistent with the typical values of $\approx 0.01-0.1$ estimated for observed gas in diverse environments and averaging scales (e.g., García-Burillo et al., 2012; Utomo et al., 2018; Sun et al., 2023). As noted in the previous section, the definitions of star formation efficiency and scales on which it is measured vary among studies. The comparison here is therefore ballpark by necessity and the trends with scale or surface density for different $\epsilon_{\mathrm{ff}}$ definitions may differ.
@@
Estimates of $\epsilon_{\mathrm{ff}}$  from the star-forming sub-pc cores to giant molecular clouds on tens of pc scales are $\approx 0.01-0.1$ (e.g., Krumholz & Tan, 2007; Evans et al., 2014; Pokhrel et al., 2021; Hu et al., 2022). Similar values of $\epsilon_{\mathrm{ff}}$ are estimated on the scales of hundreds of pc in other galaxies with a diverse range of metallicities and surface densities (e.g., Leroy et al., 2008; Utomo et al., 2018; Teng et al., 2024). Our results indicate that these characteristic values of $\epsilon_{\mathrm{ff}}$ and their universality can be plausibly explained by the turbulence-driven and feedback-regulated properties of star-forming regions.
@@
We note, however, that comparing detailed distributions of $\epsilon_{\mathrm{ff}}$ estimated in observations and simulations is not straightforward. In simulations, instantaneous values of $\epsilon_{\mathrm{ff}}$ are considered, while in observations, these values are estimated using instantaneous gas properties and star formation rate indicators that are sensitive to star formation over a certain time scale. The gas density is expected to evolve strongly as the star formation and feedback proceed in a given region (Feldmann & Gnedin, 2011). The large apparent scatter of observational estimates of $\epsilon_{\mathrm{ff}}$ thus largely reflects diverse evolutionary stages of the regions in which these estimates are made (e.g., Feldmann & Gnedin, 2011; Lee et al., 2016; Grisdale et al., 2019). Nevertheless, the simulations shed light on why the characteristic average values of $\epsilon_{\mathrm{ff}}$ do not vary with environment properties.